If M<1 is four times its supplement, find M<1 and its supplement.
I've tried to set the equation up like this:
m<1=4x-180 but i really don't know what i am doing! thank you soo much

Found 2 solutions by stanbon, bucky:
Answer by stanbon(75887)

If M<1 is four times its supplement, find M<1 and its supplement.
----------------
Let the angle (M<1) be "x".
Its supplement is 180-x
-----------------------
EQUATION:
x = 4(180-x)
x = 720-4x
5x = 720
x = 144 (the angle you call M<1)
180-144=36 ( the supplement of M<1)
============================
Cheers,
Stan H.

Answer by bucky(2189)

If you are given the measure of an angle 1, then what is the measure of the angle of its
supplement?
.
Since the supplement of an angle is the angle that must be added to make the total 180 degrees.
.
So the supplement of angle 1 (call it angle S) can be found by subtracting
the measure
of angle 1 from 180 degrees. In equation form this could be:
.
mS = 180 - m1 [call this equation 1]
.
or in words, the measure of the supplementary angle S equals 180 degrees minus the measure of
the angle 1.
.
Next you are told that the measure of angle 1 equals 4 times the measure of angle S.
In equation form this is:
.
m1 = 4*(mS) [call this equation 2]
.
But from equation 1 we know mS = 180 - m1 . Therefore for mS we can substitute 180 - m1
into equation 2 to get:
.
m1 = 4*(180 - m1)
.
Multiply out the right side. 4 times 180 = 720 and 4 times m1 = 4*m1. This makes the
equation become:
.
m1 = 720 - 4*m1)
.
You can get rid of the 4*m1 on the right side by adding 4*m1 to both sides to get:
.
5*m1 = 720
.
Solve for m1 by dividing both sides by 5 to get:
.
m1 = 720/5 = 144 degrees.
.
Since m1 = 144 degrees, its supplement is the angle that you add to 144 to get 180 degrees.
Therefore the supplement is 180 degrees minus 144 degrees which is 36 degrees.
.
Check by finding out if the m1 equals 4 times its supplement. In other words, does
.
144 degrees = 4 times 36 degrees
.
Multiply out the right side and it becomes 144 degrees, so our answer checks.
.
Hope this helps you to understand the problem and figure out what you are doing.
.